OpenStudy (anonymous):

a satellite launch rocket has a cylindrical fuel tank the fuel tank can hold the cubic meters of fuel in the tank measures the meters across what is the height of the tank in meters

OpenStudy (barreraa):

The volume of a cylinder with flat ends is (pi) x (radius squared) x (length) . Radius of the tank = 1/2 diameter = d/2. Volume = (pi) x (d/2)² x (height) Height = ( V ) / ( pi x d²/4 ) Height = ( 4V ) / ( pi d² )

OpenStudy (barreraa):

is that your answer @rarendol

OpenStudy (anonymous):

Yea thanks